OpenStudy (anonymous):

Evaluate the following integral. y=0 to y=5 and x=5y to x=0, cos(x^2)dxdy

OpenStudy (anonymous):

\[\int\limits_{0}^{5}\int\limits_{5y}^{0} f dx dy\]

OpenStudy (dumbcow):

the problem is the cos(x^2) ...there is no anti-derivative so im thinking there needs to be a change of limits but not sure what

OpenStudy (anonymous):

you can solve for y=f(x)=cos(x^2)

OpenStudy (anonymous):

yep, i think that x=0 to x=25 and y= x/5 to 5